OpenStudy (anonymous):

Mount Everest is over 2.9 × 10 4 feet tall. In standard form, that is _____. 290,000 0.00029 29,000 2,900

OpenStudy (anonymous):

1). The highest elevation on earth is Mount Everest in Asia. It is approximately 2.9 x 10^4 feet high. If a climber only made it half way up the mountain, how high did he climb? Write the answer in scientific notation. He climbed 1.45 x 10^4 feet. 2). The world population in the year 2011 was approximately 7 x 10^9 people. If the population doubles in 100 years, how many people will be on the earth in the year 2111? Write your answer in scientific notation. There will be 1.4 x 10^10 people in the year 2111. 3). A baby mouse weighs approximately 0.0022 pounds at birth. Write this weight in scientific notation. 2.2 x 10^-3 pounds. 4). The average distance from the earth to the sun is approximately 93,000,000 miles. Write this distance in scientific notation. 9.3 x 10^7 miles. 5). The average distance from the Earth to the Moon is approximately 238,000 miles. How many miles did an astronaut travel if she went from the earth to the moon and back? Write the answer in scientific notation. 4.76 x 10^5 miles. 6). The projected world population for the year 2025 is 8,000,000,000 people. Write this number in short word form. Eight billion people. 7). The world population in 1995 was approximately 5.7 billion people. Write this number in standard form. 5,700,000,000. 8). It takes light to roughly 3.0 x 10^-9 second to travel one meter. Write this time in standard form. 0.000,000,003 seconds 9). Think about a career that interests you in which you would work with very large or very small numbers. Write a paragraph about how you might use scientific notation, standard form, and short word form if you go into that career field. Provide specific examples. Scientific Notation (also called Standard Form) is a special way of writing numbers that makes it easier to use big and small numbers. Example: 4,900,000,000 1,000,000,000 = 10^9 , so 4,900,000,000 = 4.9 × 10^9 in "Scientific Notation" Example: 0.0055 would be written as 5.5 × 10^-3 Because 0.0055 = 5.5 × 0.001 = 5.5 × 10^-3 Example: 0.00012 seconds would be written as 120 × 10^-6 s, or 120 microseconds http://www.mathsisfun.com/numbers/scient...